OpenStudy (anonymous):

Dizzy question: |3x-7|=|5x+3| how do I solve this?

OpenStudy (anonymous):

solve two equations, that is it first solve \[3x-7=5x+3\] then solve \[-3x+7=5x+3\] those will give your two answers

OpenStudy (anonymous):

ok but how about the right side... do we have to change its sign too ? ... if not , why?

OpenStudy (anonymous):

well you can if you like, and you will get 4 equations to solve. but then you will see that you really get the same two equations twice

OpenStudy (anonymous):

what you have is \(|a|=|b|\) so \(a=b\) or \(a=-b\) if you write down \[a=b\] \[a=-b\] \[-a-b\] \[-a=-b\] you see that the bottom two are identical to the top two

OpenStudy (anonymous):

actually the word is "equavalent" bottom two are equivalent to top two

OpenStudy (anonymous):

ahh .. I see, Thanks a lot, :) , I always had in absolute value only one side of the equation, never twice. And that's why I had some doubts lol. Thank you very much @satellite73

OpenStudy (anonymous):

never both of them :$ ... (My english sucks !)

OpenStudy (anonymous):

\[\Large 3x-7=5x+3\]\[\Large 3x-5x=3+7\]\[\Large -2x=10\]\[\Large x=-5 \] \[-3x+7=5x+3\]\[-8x=-4\]\[x=\frac12\] lol... :P